# Phase Separation in Kitaev Chain

### v3

@@ Abstract
where $a_{j}$ is the operator for spinless fermions and $\mu$ denotes the chemical potential. $t$ and $V$ are the transfer integral and the attractive interaction between NN sites, respectively. We decouple this Hamiltonian by taking the following OPs: $p$-wave SCOP $\Delta=-i\langle a_{j+1}a_{j}\rangle$, bond OP (BOP) $\chi=\langle a^{\dagger}_{j+1}a_{j}\rangle$, and fermion density $n=\langle a^{\dagger}_{j}a_{j}\rangle$. We impose the periodic boundary condition, and assume that $\chi$ and $\Delta$ are real. The self-consistency equations can be obtained by minimizing the free energy as follows
@@
The above considerations indicate that the occurrence of PS is due to the OPs usually ignored. If we discard the term $2nV$ and retain the term $V\chi$ in $\xi_{k}$, PS would not occur. This means that the former term, which is proportoinal to the particle density in the dispersion relation $\xi_{k}$, is the primary cause of PS. On the contrary, if we drop the term $V\chi$ and keep the term $2nV$, PS may occur but the value of $V$ necessary to induce PS becomes larger. This is because the inclusion of $V\chi$ in $\xi_{k}$ reduces the band width, and so the density of states at the fermi level is increased, leading to the situation favorable to the phase transition. Then dropping the term $V\chi$ works to suppress PS.

### v4

@@ Abstract
where $a_{j}$ is the operator for spinless fermions and $\mu$ denotes the chemical potential. $t$ and $V$ are the transfer integral and the attractive interaction between NN sites, respectively. We decouple this Hamiltonian by taking the following OPs: $p$-wave SCOP $\Delta=-i\langle a_{j+1}a_{j}\rangle$, bond OP (BOP) $\chi=\langle a^{\dagger}_{j+1}a_{j}\rangle$, and fermion density $n=\langle a^{\dagger}_{j}a_{j}\rangle$. (Here we denote $n$ as a kind of OP to signal phase separation, although it is always finite.) We impose the periodic boundary condition, and assume that $\chi$ and $\Delta$ are real. The self-consistency equations can be obtained by minimizing the free energy as follows
@@
The above considerations indicate that the occurrence of PS is due to the OPs usually ignored. If we discard the term $2nV$ and retain the term $V\chi$ in $\xi_{k}$, PS would not occur. This means that the former term, which is proportional to the particle density in the dispersion relation $\xi_{k}$, is the primary cause of PS. On the contrary, if we drop the term $V\chi$ and keep the term $2nV$, PS may occur but the value of $V$ necessary to induce PS becomes larger. This is because the inclusion of $V\chi$ in $\xi_{k}$ reduces the band width, and so the density of states at the Fermi level is increased, leading to the situation favorable to the phase transition. Then dropping the term $V\chi$ works to suppress PS.